Another ttest question with the pre and post test copmpasite scores

pre

49
67
61
61
64
58
58
61
49
67
55
55
58
67
52
43
58
58
70
52
64
61
58

Post

55
70
64
67
73
88
67
82
82
73
76
88
55
88
58
76
76
76
66
76
76
61
82
Answer #1

Solution:-

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: ud = 0

Alternative hypothesis: ud ≠ 0

Note that these hypotheses constitute a two-tailed test. The null hypothesis will be rejected if the difference between sample means is too big or if it is too small.

Formulate an analysis plan. For this analysis, the significance level is 0.05. Using sample data, we will conduct a matched-pairs t-test of the null hypothesis.

Analyze sample data. Using sample data, we compute the standard deviation of the differences (s), the standard error (SE) of the mean difference, the degrees of freedom (DF), and the t statistic test statistic (t).

Before After d = A - B (d - dbar)^2

49

55 -6 68.96219282
67 70 -3 127.7882798
61 64 -3 127.7882798
61 67 -6 68.96219282
64 73 -9 28.13610586
58 88 -30 246.3534972
58 67 -9 28.13610586
61 82 -21 44.83175803
49 82 -33 349.5274102
67 73 -6 68.96219282
55 76 -21 44.83175803
55 88 -33 349.5274102
58 55 3 299.4404537
67 88 -21 44.83175803
52 58 -6 68.96219282
43 76 -33 349.5274102
58 76 -18 13.65784499
58 76 -18 13.65784499
70 66 4 335.0491493
52 76 -24 94.00567108
64 76 -12 5.310018904
61 61 0 204.6143667
58 82 -24 94.00567108
Sum 1346 1675 -329 3076.869565
Mean 58.521739 72.82608696 -14.304348 133.7769376

s = sqrt [ (\sum (di - d)2 / (n - 1) ]

s = 11.826

SE = s / sqrt(n)

S.E = 2.4659

DF = n - 1 = 23 -1

D.F = 22

t = [ (x1 - x2) - D ] / SE

t = - 5.80

where di is the observed difference for pair i, d is mean difference between sample pairs, D is the hypothesized mean difference between population pairs, and n is the number of pairs.

Since we have a two-tailed test, the P-value is the probability that a t statistic having 22 degrees of freedom is more extreme than - 5.80; that is, less than - 5.80 or greater than 5.80.

Thus, the P-value = less than 0.0001

Interpret results. Since the P-value (0.00) is less than the significance level (0.05), we have to reject the null hypothesis.

Reject H0. The mean difference appears to differ from zero.